OpenStudy (anonymous):

Which equation has a solution of n = -1

OpenStudy (anonymous):

The one you just wrote....

OpenStudy (anonymous):

2n+2=n+1

OpenStudy (anonymous):

1/2 * n - 6 = 7 -6*n+8 = -2 2*(n + 8) = -2 2*n + 12 -3*n=15

OpenStudy (anonymous):

n+1 = 0

OpenStudy (anonymous):

is it the first one?

OpenStudy (anonymous):

what about this=2n+2=n+1

OpenStudy (anonymous):

what about n+1 = 0? (out of an infinite number of such equations..)

OpenStudy (anonymous):

it's not an option, so it's wrong.

OpenStudy (anonymous):

Where are the options, I see no options....

OpenStudy (anonymous):

the ones i wrote....

OpenStudy (anonymous):

a. 1/2 * n - 6 = 7 b. -6*n+8 = -2 c. 2*(n + 8) = -2 d.2*n + 12 -3*n=15

OpenStudy (anonymous):

A

OpenStudy (anonymous):

U can just put -1 into each one till you see the one where it works....